Monochromatic light λ=520nm is aimed through a pinhole, producing a diffraction pattern on a wall 1.75m away. If the radius of the second dark ring 2.50mm, what’s the pinhole’s diameter?
Answer #1


use in diffraction Y = mLR/D

Y = 2 * 520 e-9 * 1.75/(2.5 e-3)

Y = 0.728 mm

diameter 2Y = 2* 0.728

2Y = 1.456 mm